OpenStudy (anonymous):

P=258e^(0.024t) When will the population reach 2512 people?

OpenStudy (anonymous):

you miss a number

OpenStudy (anonymous):

t = 94.82 years

OpenStudy (anonymous):

how did you get that @smr_kml ?

OpenStudy (anonymous):

it's the same steps but he wrote 25 instead of 258

OpenStudy (anonymous):

oh shnap yea sorry. 2512 = 258e^(0.024t) 9.736 = e^(0.024t) ln 9.736 = 0.024t t = 94.828 years

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

Do you get it @Aprilmbug ?

OpenStudy (anonymous):

Okay thank you!!! I kept missing a step.

OpenStudy (anonymous):

you just have to take the natural log of 2512/258 and then you will got 2.27 = 0.024t

OpenStudy (anonymous):

and then keep the t on one side t = 2.27/0.024= 94.82 years

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

Yeah did a similar problem on my online class and got it right. Thanks!

OpenStudy (anonymous):

you welcome anytime